# Elastic Decision Transformer

### v1

@@ 3 Elastic Decision Transformer
We first provide essential background knowledge (Sec. 3.1) and discuss the motivation behind our approach (Sec. 3.2). Subsequently, we present our novel objective designed for training the EDT, which integrates expectile regression (Sec. 3.3). We also detail the action inference process employed during testing (Sec. 3.4).
@@ 3.3 Training objective for Maximum In-Support Return
where $\mathcal{L}_{\text{observation}}$ and $\mathcal{L}_{\text{action}}$ are computed with a mean square error, $\mathcal{L}_{\text{return}}$ is a cross-entropy loss, and $\mathcal{L}_{\text{max}}$ is an empirical estimate of Equation 2. We set $c_{r}=0.001$ to balance scale differences between mean square error and cross-entropy loss. In tasks with discrete action spaces like Atari, we optimize the action space as the return objective $\mathcal{L}_{\text{return}}$ using cross-entropy with weight $10c_{r}$. Our training method extends the work of lee2022multi by estimating the maximum expected return value for a trajectory using Equation 2. This estimation aids in comparing expected returns of different trajectories over various history lengths. Our proposed method is not only easy to optimize, but can also be conveniently integrated with other DT variants. As such, it marks a significant advance in developing efficient offline reinforcement learning approaches for complex decision-making tasks.
@@ 3.4 Action Inference During Test time
To sample from expert return distribution $P(R^{t},...|\text{expert}^{t})$, we adopt an approach similar to lee2022multi by applying Bayes’ rule $P(R^{t},...|\text{expert}^{t})\propto P(\text{expert}^{t}|R^{t},...)P(R^{t},...)$ and approximate the distribution of expert-level return with inverse temperature $\kappa$ [^1] We set $\kappa$ to $10$ in all our experiments. (kappen2012optimal, ; toussaint2009robot, ; todorov2006linearly, ; shachter1988probabilistic, ):
@@
Task DT-1 IQL-1 EDT-1 (Ours) hopper $51.2$ $59.8$ $\bm{76.9}$ walker $29.8$ $52.6$ $\bm{74.1}$ halfcheetah $30.5$ 40.4 $36.8$ ant $79.8$ $82.3$ $\bm{88.6}$ sum $191.3$ $235.1$ 276.4 Table 2: Evaluation results on multi-task regime. Mean of 5 random training initialization seeds, 100 evaluations each. The training dataset is a mixture of medium-replay datasets from the four locomotion tasks. Our main result is highlighted. Refer to caption Figure 5: The average HNS comparison on $20$ Atari games. The results are evaluated with three trials.
@@ 4.1 Baseline Methods
It is important to note that QDT and TS+BC are specifically designed to achieve trajectory stitching. QDT accomplishes this by substituting collected return values with estimates derived from Conservative Q-Learning (kumar2020conservative, ). Conversely, TS+BC employs a model-based data augmentation strategy to bring about the stitching of trajectories.
@@ 4.2 Single-Task Offline Reinforcement Learning
Dataset EDT (w=20) EDT (w=5) EDT (Ours) hopper-m $63.6\pm 5.2$ $57.8\pm 7.0$ $\bm{63.5\pm 5.8}$ hopper-mr $67.6\pm 27.7$ $62.6\pm 26.9$ $\bm{89.0\pm 8.3}$ walker-m $65.6\pm 11.7$ $62.6\pm 14.3$ $\bm{72.8\pm 6.2}$ walker-mr $64.5\pm 12.9$ $44.3\pm 8.7$ $\bm{74.8\pm 4.9}$ halfcheetah-m $42.0\pm 0.4$ $42.3\pm 0.8$ $\bm{42.5\pm 0.9}$ halfcheetah-mr $33.5\pm 8.0$ $36.4\pm 7.4$ $\bm{37.8\pm 1.5}$ ant-m $90.8\pm 16.0$ $95.6\pm 8.0$ $\bm{97.9\pm 8.0}$ ant-mr $80.0\pm 17.0$ $82.3\pm 15.9$ $\bm{92.0\pm 4.1}$ sum $507.6$ $483.9$ 570.3 Table 3: Mean of 5 random training initialization seeds, 100 evaluations each. In the Dataset column, the "m" indicates medium dataset, "mr" indicates medium-replay dataset. The "w" stands for history length. Our main results are highlighted. Refer to caption Figure 6: Ablation study on expectile level $\alpha$. Expectile objective reduces to Mean Square Error when $\alpha=0.5$. Evaluated on Hopper and medium-replay dataset.
We conclude our locomotion results in Table 1. To simplify implementation, we train a casual transformer decoder from scratch instead of adopting the pre-trained GPT-2 model as zheng2022online ; chen2021decision . In addition, since the proposed model estimates the return of the current sequence, reward information is not required during test time.
@@ 4.3 Multi-Task Offline Reinforcement Learning
Locomotion. In the locomotion multi-task experiment, we maintain the same model architecture as in the single-task setting. By confining the dataset to medium-replay datasets from four tasks, we increase task complexity and necessitate the offline RL approach to learn and execute these tasks concurrently, while effectively utilizing trajectories generated by random policies. As depicted in Figure 2, our proposed EDT successfully accomplishes all four tasks simultaneously without much performance compromise.
@@
Our experimental results align with those of lee2022multi , highlighting that Q-learning-based offline RL approaches encounter difficulties in learning a multi-task policy on Atari games. Despite IQL achieving the highest score in Table1, it demonstrates relative inadequacy in simultaneous multi-task learning as indicated in Table 2 and Figure 5. We leave the raw scores of the $20$ Atari games in Appendix B.
@@
To validate the benefits of this dynamic history length approach, we tested the EDT using both fixed and variable history lengths. The results, summarized in Table 3, show that the variable history length outperforms the fixed ones, particularly on the "medium-replay" datasets.
@@ 4.5 Ablation Study on Expectile Level $\alpha$
The results, as displayed in Figure 6, suggest that when expectile regression is able to accurately approximate the maximizer, specifically at higher expectile levels, we observe both a higher average performance and lower standard deviation. This suggests that accurate selection of history length not only stabilizes performance but also enhances scores. Conversely, as the expectile level approaches $0.5$, the expectile regression’s objective shifts towards a mean square error, leading to an estimated value that is more of a mean value than a maximal one. This change makes it a less effective indicator for optimal history length. As a result, we can see a deterioration in EDT’s score as the expectile level drops too low, and an increase in standard deviation, indicating inconsistency in the selection of an effective history length.
@@ 5 Related Work
Decision Transformer and its variants (chen2021decision, ; lee2022multi, ; yamagata2022q, ) have been a promising direction for solving offline RL from the perspective of sequence modeling. Trajectory Transformer (TT) janner2021offline models distributions over trajectories using transformer architecture. The approach also incorporates beam search as a planning algorithm and demonstrates exceptional flexibility across various applications, such as long-horizon dynamics prediction, imitation learning, goal-conditioned reinforcement learning, and offline reinforcement learning.
@@ A.1 Step Size Ablation
Figure 8 shows that a narrowed search space leads to a decline in return performance and an increase in standard deviation, corroborating results from Table 3. EDT with a restricted history search space behaves more like EDT with a fixed history length. We also found the return to be optimized when the history length is set to $4$, possibly due to challenges in estimating the return in offline settings [26, 28], where direct environment interactions are prohibited. By increasing the step size appropriately, the return inference becomes more stable and resilient to noise and disruptions. Lastly, the inference speed when $\delta=4$ is substantially faster (by a factor of four) than when $\delta=1$.
@@ Appendix B Experiment Details
B.1 Learning on Locomotion Tasks
In Section 4, we discussed the application of our approach to a multi-task learning scenario. Specifically, we consolidated medium-replay datasets from Ant, HalfCheetah, Hopper, and Walker2d. To standardize returns across these varied environments, we used scaling statistics (maximum return, minimum return, and return scale) from the official Decision Transformer repository (https://github.com/kzl/decision-transformer). As detailed in Section 3.3, we further segmented the return into 120 discrete buckets and, following the approach of [20, 30], sampled from the top 85th percentile logits within the discrete return distribution.
However, it’s important to highlight that our return maximizer, $\tilde{R}$, estimates the scaled value directly rather than the discretized one. For the sequence $\langle...,\mathbf{o},R,a,...\rangle$, we supplement each token embedding with a learned positional embedding. Given the differing state dimensions across tasks, we employ an additional state embedding layer. This layer transforms the raw state representation into a consistent state embedding size, while keeping the remainder of the architecture unchanged across tasks.
@@ B.1 Learning on Locomotion Tasks
History Length Search Heuristic. In Algorithm 1, we illustrated that a larger $\delta$ value allows us to infer actions more rapidly. Adopting this method needs to balance inference speed with search accuracy. Based on the concept introduced in Sec. 3.2, the optimal history length at the current state $s_{t+1}$ might be close to that of the previous state $s_{t}$. Therefore, given the optimal length $l_{t}$ at time step $t$, we search within the range $\{l_{t}-\Delta,l_{t}-\Delta+1...,l_{t}+\Delta\}$ for the optimal length $l_{t+1}$ at the next step. Our results show that with $\Delta=3$, we can achieve an inference speed that is around three times faster than Algorithm 1 with $\delta=1$, with slight improvement in performance and lower variance.
B.2 Multi-Task Learning on Atari Games
@@
Given the complexity and potential pitfalls of learning on 20 Atari games from scratch, we use GPT-v2 to initialize the transformer. This step improves both the convergence rate and the quality of the learned policy. Our dataset consists of 2 training runs from [2], with each run featuring rollouts from 50 checkpoints. Lastly, we enhance the dataset with image augmentations, including random cropping and rotation.
B.3 Atari Games
@@ Appendix D Architecture Details
We adopt the casual transformer decoder architecture and process the sequences as follows.
@@
Return maximizer loss coefficient ($\mathcal{L}_{\text{max}}$) 1 1
Number of return output bins 120 120

### v6

@@ 3.3 Training objective for Maximum In-Support Return
where $\mathcal{L}_{\text{observation}}$ and $\mathcal{L}_{\text{action}}$ are computed with a mean square error, $\mathcal{L}_{\text{return}}$ is a cross-entropy loss, and $\mathcal{L}_{\text{max}}$ is an empirical estimate of Equation 2. We set $c_{r}=0.001$ to balance scale differences between mean square error and cross-entropy loss. In tasks with discrete action spaces like Atari, we optimize the action space as the tokenized return objective $\mathcal{L}_{\text{return}}$ using cross-entropy with weight $10c_{r}$. Our training method extends the work of lee2022multi by estimating the maximum expected return value for a trajectory using Equation 2. This estimation aids in comparing expected returns of different trajectories over various history lengths. Our proposed method is not only easy to optimize, but can also be conveniently integrated with other DT variants. As such, it marks a significant advance in developing efficient offline reinforcement learning approaches for complex decision-making tasks.
@@ 3.4 Action Inference During Test time
To sample from expert return distribution $P(R^{t},...|\text{expert}^{t})$, we adopt an approach similar to lee2022multi by applying Bayes’ rule $P(R^{t},...|\text{expert}^{t})\propto P(\text{expert}^{t}|R^{t},...)P(R^{t},...)$ and approximate the distribution of expert-level return with inverse temperature $\kappa$ (kappen2012optimal, ; toussaint2009robot, ; todorov2006linearly, ; shachter1988probabilistic, ):
@@ 4.1 Baseline Methods
It is important to note that QDT and TS+BC are specifically designed to achieve trajectory stitching. QDT accomplishes this by substituting collected return values with estimates derived from Conservative Q-Learning (kumar2020conservative, ). Conversely, TS+BC employs a model-based data augmentation strategy to bring about the stitching of trajectories."
@@ 4.2 Single-Task Offline Reinforcement Learning
We conclude our locomotion results in Table 1. Since the proposed model estimates the return of the current sequence, reward information is not required during test time.
@@ 4.3 Multi-Task Offline Reinforcement Learning
Locomotion. In the locomotion multi-task experiment, we maintain the same model architecture as in the single-task setting. By confining the dataset to medium-replay datasets from four tasks, we increase task complexity and necessitate the offline RL approach to learn and execute these tasks concurrently, while effectively utilizing trajectories generated by random policies. As depicted in Table LABEL:tab:edt_all_mujoco, our proposed EDT successfully accomplishes all four tasks simultaneously without much performance compromise.
@@
Our experimental results align with those of lee2022multi , highlighting that Q-learning-based offline RL approaches encounter difficulties in learning a multi-task policy on Atari games. Despite IQL achieving the highest score in Table1, it demonstrates relative inadequacy in simultaneous multi-task learning as indicated in Table LABEL:tab:edt_all_mujoco and Figure LABEL:fig:atari. We leave the raw scores of the $20$ Atari games in Appendix B.
@@
To validate the benefits of this dynamic history length approach, we tested the EDT using both fixed and variable history lengths. The results, summarized in Table LABEL:tab:edt_length, show that the variable history length outperforms the fixed ones, particularly on the "medium-replay" datasets.
@@ 4.5 Ablation Study on Expectile Level $\alpha$
The results, as displayed in Figure LABEL:fig:exp_level, suggest that when expectile regression is able to accurately approximate the maximizer, specifically at higher expectile levels, we observe both a higher average performance and lower standard deviation. This suggests that accurate selection of history length not only stabilizes performance but also enhances scores. Conversely, as the expectile level approaches $0.5$, the expectile regression’s objective shifts towards a mean square error, leading to an estimated value that is more of a mean value than a maximal one. This change makes it a less effective indicator for optimal history length. As a result, we can see a deterioration in EDT’s score as the expectile level drops too low, and an increase in standard deviation, indicating inconsistency in the selection of an effective history length.
@@ 5 Related Work
Decision Transformer and its variants (chen2021decision, ; lee2022multi, ; yamagata2022q, ; xu2022prompting, ; xu2023hyper, ) have been a promising direction for solving offline RL from the perspective of sequence modeling. Trajectory Transformer (TT) janner2021offline models distributions over trajectories using transformer architecture. The approach also incorporates beam search as a planning algorithm and demonstrates exceptional flexibility across various applications, such as long-horizon dynamics prediction, imitation learning, goal-conditioned reinforcement learning, and offline reinforcement learning.
@@ A.1 Step Size Ablation
Figure 6 shows that a narrowed search space leads to a decline in return performance and an increase in standard deviation, corroborating results from Table LABEL:tab:edt_length. EDT with a restricted history search space behaves more like EDT with a fixed history length. We found that the EDT works best when we chose a search step of $4$. This could be because it is tough to make good estimates without being able to interact with the environment directly, as mentioned in previous studies [26, 28]. By increasing the step size, we were able to make our estimates more reliable and less affected by noise and other problems.
@@ Appendix B Experiment Details
B.1 Loss Computation Details
In this section, we detail the computation of the tokenized return loss $\mathcal{L}_{\text{return}}$ and the observation loss $\mathcal{L}_{\text{observation}}$ in Eq 1. For $\mathcal{L}_{\text{return}}$, we first tokenize the predicted return and the target return into $n$ bins bounded by the MAX and MIN scores obtained from D4RL [15]. We then compute the cross-entropy loss between the predicted value and the target value:
$\mathcal{L}_{\text{return}}=CE(\hat{R}^{t},R^{\prime t}),$ (5)
where $R^{\prime t}$ suggests the predicted tokenized return. For $\mathcal{L}_{\text{observation}}$, we predict the next observation by minimizing the mean square error.
B.2 Learning on Locomotion Tasks
In Section 4, we discussed the application of our approach to a multi-task learning scenario. Specifically, we consolidated medium-replay datasets from Ant, HalfCheetah, Hopper, and Walker2d. To standardize returns across these varied environments, we used scaling statistics (maximum return, minimum return, and return scale) from the official Decision Transformer repository (https://github.com/kzl/decision-transformer). As detailed in Section 3.3, we further segmented the return into $60$ discrete buckets and, following the approach of [20, 30], sampled from the top 85th percentile logits within the discrete return distribution.
However, it’s important to highlight that our return maximizer, $\tilde{R}$, estimates the scaled value directly rather than the discretized one. For the sequence $\langle...,\mathbf{o},R,a,...\rangle$, we supplement each token embedding with a learned positional embedding. Given the differing state and action dimensions across tasks, we employ an additional state and action embedding layer. This layer transforms the raw state and action representation into a consistent state and action embedding size, while keeping the remainder of the architecture unchanged across tasks.
@@ B.2 Learning on Locomotion Tasks
History Length Search Heuristic. In Algorithm 1, we illustrated that a larger $\delta$ value allows us to infer actions more rapidly. Adopting this method needs to balance inference speed with search accuracy. Based on the concept introduced in Sec. 3.2, the optimal history length at the current state $s_{t+1}$ might be close to that of the previous state $s_{t}$. Therefore, given the optimal length $l_{t}$ at time step $t$, we search within the range $\{l_{t}-\Delta,l_{t}-\Delta+1...,l_{t}+\Delta\}$ for the optimal length $l_{t+1}$ at the next step.
B.3 Multi-Task Learning on Atari Games
@@
Given the complexity and potential pitfalls of learning on 20 Atari games from scratch, we use a subset of GPT-2 to initialize the transformer decoder. This step improves both the convergence rate and the quality of the learned policy. Our dataset consists of 2 training runs from [2], with each run featuring rollouts from 50 checkpoints. Lastly, we enhance the dataset with image augmentations, including random cropping and rotation.
B.4 Atari Games
@@ Appendix D Architecture Details
We adopt the causal transformer decoder architecture and process the sequences as follows.
@@
Return maximizer loss coefficient ($\mathcal{L}_{\text{max}}$) 0.5 0.5
Number of return output bins 60 120